# MLLM-based Discovery of Intrinsic Coordinates and Governing Equations from High-Dimensional Data

### v1

@@ Abstract
Discovering governing equations from scientific data is crucial for understanding the evolution of systems, and is typically framed as a search problem within a candidate equation space. However, the high-dimensional nature of dynamical systems leads to an exponentially expanding equation space, making the search process extremely challenging. The visual perception and pre-trained scientific knowledge of multimodal large language models (MLLM) hold promise for providing effective navigation in high-dimensional equation spaces. In this paper, we propose a zero-shot method based on MLLM for automatically discovering physical coordinates and governing equations from high-dimensional data. Specifically, we design a series of enhanced visual prompts for MLLM to enhance its spatial perception. In addition, MLLM’s domain knowledge is employed to navigate the search process within the equation space. Quantitative and qualitative evaluations on two representative types of systems demonstrate that the proposed method effectively discovers the physical coordinates and equations from both simulated and real experimental data, with long-term extrapolation accuracy improved by approximately 26.96% compared to the baseline [^1] Code will be available after review..
1 Introduction
@@
Nonlinear dynamics are prevalent across numerous fields in science and engineering, such as astrophysics, fluid mechanics, and materials science [1, 2, 3]. Identifying the symbolic equations governing the dynamics from data is crucial for understanding and predicting such systems, a task known as symbolic regression [4, 5, 6]. Traditional symbolic regression approaches typically assume that the physical coordinates are of low dimensionality and focus on developing more efficient and accurate search algorithms [7, 8, 9]. However, real-world scientific scenarios are often hindered by the high-dimensional nature of observational data, making it difficult to navigate the exponentially expanding symbolic space, even though a significant portion of the high-dimensional variables may be redundant [10, 11]. Reasoning physical knowledge from high-dimensional observational data to facilitate predictive understanding [12, 13] poses a novel and interdisciplinary challenge.
@@
From Kepler’s discovery of planetary motion laws to the most advanced symbolic regression techniques, the discovery of equations has historically relied on scientists manually processing observational data to obtain compact physical quantities in suitable coordinate systems [10, 14]. However, automatically discovering both physical coordinates and the governing equations on them from high-dimensional data is a challenging problem. With advancements in deep learning, especially visual models [15, 16, 17], discovering physical laws from high-dimensional data, like videos, has become more feasible. Unsupervised motion video prediction with neural networks integrated with parameterized physical modules (e.g., neural ODE [18]) allows for implicit learning of dynamics from high-dimensional observations [19, 20, 21, 22, 23]. Another approach is to embed physical engines as inductive biases into the model pipeline, compelling the neural network to infer physical quantities that satisfy the governing physical laws from video streams [24, 25, 26, 27, 28]. However, these methods separate the discovery of coordinates and equations, presuming known dynamics equations or physics engines before identifying physical quantities. Recent efforts have succeeded in jointly discovering physical coordinates and equations [29, 14, 30], yet they still rely on prior knowledge of equation forms and state dimensions, limiting their generalization to unfamiliar systems.
@@
High-dimensional data challenges traditional optimization methods like Bayesian optimization due to the lack of well-defined quantitative evaluations for intrinsic coordinates and governing equations. Large language models (LLMs) excel in problem-solving by leveraging pretrained knowledge and enhancing prompts for scientific discovery across various domains. [31, 32, 33]. Multimodal large language models (MLLMs) with visual encoders achieve spatial understanding akin to human visual intelligence. [34]. Recent studies [12] show MLLMs can deduce physical properties and motion patterns from images without prior knowledge, revealing their potential in high-dimensional data. Moreover, most symbolic regression methods, based on genetic programming, search equation spaces through mutation and recombination. By leveraging the pretraining knowledge of LLM to guide this search, equation discovery can be more efficient and avoid suboptimal solutions [35, 36]. Thus, we can anticipate that a well-designed reasoning framework and visual prompts could harness MLLMs for visual perception and scientific discovery, enabling automatic identification of joint physical coordinates and governing equations.
@@
2 Problem Fomulation
In the task of inferring governing equations for high-dimensional data, the goal is to find a compact and accurate symbolic expression. Here, we consider a class of video-like systems governed by low-dimensional coordinates, that is, systems whose dynamics evolve on a low-dimensional manifold [29]. Given the dataset $\mathcal{D}=\{x_{i}\}^{n}_{i=1}$, where $x_{i}\in\mathbb{R}^{D}$ is the $i$-th observed state (i.e., a frame of video), we aim to uncover the underlying mathematical relationships such that $\frac{dx}{dt}=f(x)$. This fundamentally relies on an effective coordinate system $\mathcal{Z}$, in which the dynamics have a simple low-dimensional representation $\frac{dz}{dt}=\hat{f}(z)$. Therefore, this type of problem consists of two coupled subproblems:
@@ 3 Methodology
3.1 Discovering physical coordinates
We consider two types of high-dimensional systems. The first type is the pixel coordinate system, where physical laws are presented in the form of Cartesian coordinates of rigidly moving objects [14, 30, 12]. MLLM needs to identify the intrinsic physical coordinate system by locating the trajectory of the target object’s motion through a sequence of coordinates. The second type is the more complex latent coordinate system, whose spatiotemporal evolution is governed by latent low-dimensional variables, presenting periodic or quasiperiodic oscillatory patterns [29, 37]. We sample a small number of frames from the data and ask MLLM to determine their type in order to initiate different locating tools.
3.1.1 Detection for Pixel Coordinate
@@
Plot Replayer displays MLLM’s initial locating results as markers (e.g., red dots) on the original frame and calls MLLM to compare the visual discrepancy between the target object and the detection location for error correction. As a feedback iteration step in the reasoning chain, the replayer is used after the initial locating of each frame. However, for the sake of simplicity in the formula, we have incorporated it into the $\mathrm{LocateTool}$ functionality in Algorithm 1.
After serially detecting the state of each frame, we introduce a filter tool to smooth the physical coordinate trajectory in order to enhance the continuity of motion between frames. Specifically, we designed a feedback correction module, $\mathrm{Filter}$, based on the Savitzky-Golay filter [38]. We plot the recognized trajectory as a visual prompt and ask MLLM to determine the window length $h$ and polynomial order $p$ based on the evolution trend and periodicity of the trajectory. Each filtering result is visualized alongside the originally recognized trajectory as the visual prompt for the next iteration, until MLLM determines that the smoothing has preserved the evolution trend while maintaining inter-frame coordinate continuity. The pseudocode for serially extracting the physical coordinate sequence from the high-dimensional motion video is shown in Algorithm 1.
3.1.2 Detection for Latent Coordinate
The task of discovering physical coordinates in the latent coordinate system is no longer visual target localization, but rather the encoding of intrinsic variables that capture complex spatiotemporal evolution patterns [10]. Here, intrinsic variables refer to the minimal set of variables that describe all the information of a dynamic system [39]. Although prior work has developed manifold learning methods such as Koopman operator theory and delayed embedding [40, 41], these models heavily rely on prior assumptions about the system’s intrinsic dimensions (i.e., the number of intrinsic variables).
@@ 3.1.2 Detection for Latent Coordinate
3.2 Reasoning governing equations
We adopt the sparse identification method for nonlinear dynamics (SINDy) [42] as the starting point for our approach. Inspired by the fact that latent governing equations typically contain only a finite number of terms, SINDy identifies the equations by optimizing the coefficients of predefined symbolic terms (e.g. $sin$, $exp$, and $x^{2}$). Specifically, SINDy defines a symbolic library $\Theta(z)=[\theta_{1}(z),\theta_{2}(z),...,\theta_{k}(z)]\in\mathbb{R}^{d\times k}$, where $\theta_{i}$ represents a candidate symbolic term. The symbolic library is linearly combined to form a deterministic system $\frac{dz}{dt}=\Theta(z)\Xi$, where $\Xi=[\xi_{1},\xi_{2},...,\xi_{k}]\in\mathbb{R}^{k\times 1}$ is the coefficient to be optimized for each symbolic term. Given the coordinate sequence $Z=\{z_{i}\}^{n}_{i=1}$ and its derivative sequence $\dot{Z}=\{\dot{z_{i}}\}^{n}_{i=1}$, SINDy minimizes the loss function
@@ 3.2 Reasoning governing equations
3.2.1 Hypothesis generation
Previous SINDy-based works have succeeded [43, 44], but rely on a high-quality predefined symbolic library, often needing prior system knowledge to manually filter out unnecessary terms. To address systems without prior knowledge, we design a hypothesis generation module to mine the physical knowledge pre-learned by MLLM and automatically select potential symbolic terms. The hypothesis generation step utilizes the frozen MLLM to propose diverse and promising candidate symbolic terms. At each iteration $t$, MLLM uses replay samples from the experience pool (Sec. 3.2.3) as text and visual prompts, suggesting the candidate symbolic library $\Theta_{t}$.
3.2.2 Hypothesis Assessment
@@
$L_{AE-SINDy}=\frac{1}{n}\sum^{n}_{i=1}||x_{i}-\psi_{d}(x_{i})||^{2}_{2}+||\nabla_{x_{i}}(z_{i})\dot{x}_{i}-\Theta_{t}(z_{i})\Xi_{t}||^{2}_{2}+\eta|\Xi|$ (2)
@@
We employ the $R^{2}$ metric as a quantitative measure of how well the equation fits the data. Additionally, we evaluate the number $length_{t}$ of symbolic terms in the equation to quantify its simplicity. In symbolic regression tasks, the goal of maintaining equation simplicity is to avoid overfitting to the limited observational data. At this stage, the candidate symbolic library hypothesis $\Theta_{t}$ is assessed based on both the $R^{2}$ metric and the equation length to determine its $fitness_{t}$. Thus, the equation discovery process is divided into two steps: (i) constructing the symbolic library, and (ii) optimizing based on stochastic gradient using PyTorch [45]. This process aligns with the philosophy of human scientific discovery [36, 46, 47], where the MLLM makes decisions on the discrete components of the SINDy symbolic library based on pretrained knowledge and prior experiences in prompts, akin to the empirical hypotheses made by human scientists during scientific discovery.
3.2.3 Experience management
@@ 4.1 Datasets
Pixel coordinate system. We validate the effectiveness of the proposed method on the pixel coordinate system using several dynamical systems shown in Figure 2, including: Linear, Cubic, Circular, Van Der Pol (VDP), Glider, and Exp equations. These equations include various characteristics such as linear terms, nonlinear terms, and significant differences in time scales [48], which together form an evaluation benchmark across different levels of difficulty. The video data is generated by simulating physical trajectories, with details provided in the Appendix A.
@@
  • Lambda–Omega (LO) equation [29] generates spiral wave structures featuring two oscillating spatial modes.
  • Brusselator (Bruss) equation [49] models the concentration dynamics in chemical reactions, where the system’s trajectory eventually reaches a limit cycle attractor.
  • Shallow-Water (Water) equation [50], derived from the compressible Navier-Stokes equations,describes fluid dynamics in shallow water scenarios, commonly applied in large-scale geophysical flows and tsunami simulations.
  • Real video [51] visualizes the formation of Kármán vortex streets as water flows past a cylindrical obstacle with a Reynolds number of 171, using dye to highlight the vortex pattern.
@@
4.2 Experimental Setup
Models. We utilize the GPT-4o [52] in our experiments to explore the performance of our proposed VER. We validate the improvements of VER over the basic GPT-4o in the ablation study. The penalty coefficient $\eta$ in Equation 1 is set to $0.01$ and we explore the performance of hyperparameter search in the appendix B.2.
Baselines. Related works [14, 53] on discovering coordinates from pixel coordinate videos typically employ a combination of visual localization models and symbolic regression models. We adopt a recent work [14], Video-SINDy, as the baseline method for the pixel coordinate system. In addition, we also replace the coordinate encoder in Video-SINDy with SAM [54] as a zero-shot baseline, SAM-SINDy. For the latent coordinate system, the most relevant work [29] constructs an end-to-end model combining an autoencoder and SINDy (AE-SINDy) to interpret low-dimensional patterns in high-dimensional data. Since it requires prior knowledge of the low-dimensional space dimensions, we ensure that its dimensions are consistent with those discovered by VER.
@@ 4.2 Experimental Setup
4.3 System with Pixel Coordinate
@@
Figure 2 shows VER’s reasoning results in the pixel coordinate system, including governing equations and inferred trajectories compared to the optimal baseline. VER uses only raw video input, without needing motion type or equation terms. VER’s trajectories closely align with ground truth over time, demonstrating MLLM’s spatial understanding through locating prompts, as supported by [34]. The inferred ODE terms perfectly match the ground truth, indicating VER’s ability to uncover intrinsic dynamics in pixel space.
@@
We assess if VER and baseline methods’ discovered equations include all correct terms, count incorrect terms, and evaluate fitting accuracy ($R^{2}$). We numerically integrate discovered equations to predict trajectories, comparing them to real ones in Table 1. VER consistently outperforms baselines, especially in long-term predictions, where inaccuracies are magnified. This highlights VER’s superior efficiency and accuracy over traditional visual symbolic regression methods. Additionally, we collect physical motion videos from public sources to confirm VER’s reliability on real systems (Appendix B.1).
@@
Table 1: Average performance of pixel coordinate systems over 10 runs with varying seeds. $R^{2}@n$ indicates $R^{2}$ for $n$ steps. Best results are in bold. $R^{2}$ scale is $10^{-1}$.
Case Method Equation Prediction
Terms Found False Positives $R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear Video-SINDy Yes $0.91\pm 0.34$ $9.46\pm 0.19$ $8.99\pm 0.31$ $8.67\pm 0.55$
SAM-SINDy Yes $0.45\pm 0.13$ $8.95\pm 0.30$ $8.72\pm 0.28$ $3.50\pm 1.31$
Cubic Video-SINDy No $3.52\pm 1.12$ $8.16\pm 0.70$ $6.79\pm 1.83$ $2.63\pm 1.92$
SAM-SINDy No $2.17\pm 1.20$ $7.64\pm 0.89$ $6.99\pm 1.22$ $2.16\pm 1.77$
Circular Video-SINDy Yes $0.31\pm 0.10$ $9.76\pm 0.11$ $9.81\pm 0.25$ $9.03\pm 0.97$
SAM-SINDy Yes $0.12\pm 0.03$ $9.21\pm 0.09$ $9.93\pm 0.13$ $9.32\pm 0.35$
VER Yes 0 $\mathbf{9.91\pm 0.07}$ $\mathbf{9.99\pm 0.00}$ $\mathbf{9.67\pm 0.45}$
VDP Video-SINDy Yes $2.31\pm 0.65$ $8.92\pm 0.81$ $8.17\pm 0.64$ $4.91\pm 1.29$
SAM-SINDy Yes $2.45\pm 0.59$ $8.82\pm 0.27$ $9.17\pm 0.39$ $6.74\pm 0.84$
VER Yes $\mathbf{0.80\pm 0.40}$ $\mathbf{9.69\pm 0.08}$ $\mathbf{9.35\pm 0.19}$ $\mathbf{7.28\pm 0.72}$
EXP Video-SINDy Yes $2.76\pm 0.88$ $9.28\pm 0.81$ $9.31\pm 0.17$ $8.98\pm 0.32$
SAM-SINDy Yes $3.21\pm 1.17$ $9.54\pm 0.16$ $9.56\pm 0.09$ $9.48\pm 0.24$
VER Yes $\mathbf{1.34\pm 0.29}$ $\mathbf{9.94\pm 0.06}$ $\mathbf{9.88\pm 0.11}$ $\mathbf{9.78\pm 0.15}$
Glider Video-SINDy No $2.18\pm 0.49$ $9.35\pm 0.20$ $9.77\pm 0.11$ $9.08\pm 0.27$
SAM-SINDy No $2.23\pm 0.56$ $8.52\pm 0.36$ $9.86\pm 0.38$ $3.16\pm 2.73$
VER Yes $\mathbf{0.84\pm 0.17}$ $\mathbf{9.78\pm 0.08}$ $\mathbf{9.90\pm 0.07}$ $\mathbf{9.23\pm 0.18}$

### v2

@@ Abstract
Discovering governing equations from scientific data is crucial for understanding the evolution of systems, and is typically framed as a search problem within a candidate equation space. However, the high-dimensional nature of dynamical systems leads to an exponentially expanding equation space, making the search process extremely challenging. The visual perception and pre-trained scientific knowledge of multimodal large language models (MLLM) hold promise for providing effective navigation in high-dimensional equation spaces. In this paper, we propose a zero-shot method based on MLLM for automatically discovering physical coordinates and governing equations from high-dimensional data. Specifically, we design a series of enhanced visual prompts for MLLM to enhance its spatial perception. In addition, MLLM’s domain knowledge is employed to navigate the search process within the equation space. Quantitative and qualitative evaluations on two representative types of systems demonstrate that the proposed method effectively discovers the physical coordinates and equations from both simulated and real experimental data, with long-term extrapolation accuracy improved by approximately 26.96% compared to the baseline.
Introduction
@@
Nonlinear dynamics are prevalent across numerous fields in science and engineering, such as astrophysics, fluid mechanics, and materials science (Tenachi, Ibata, and Diakogiannis 2023; Wang, Wagner, and Rondinelli 2019; Li et al. 2025c, b). Identifying the symbolic equations governing the dynamics from data is crucial for understanding and predicting such systems, a task known as symbolic regression (Langley 1981; Schmidt and Lipson 2009; Cornelio et al. 2023). Traditional symbolic regression approaches typically assume that the physical coordinates are of low dimensionality and focus on developing more efficient and accurate search algorithms (Biggio et al. 2021; Shojaee et al. 2023; La Cava et al. 2021). However, real-world scientific scenarios are often hindered by the high-dimensional nature of observational data, making it difficult to navigate the exponentially expanding symbolic space, even though a significant portion of the high-dimensional variables may be redundant (Chen et al. 2022; Pope et al. 2021). Reasoning physical knowledge from high-dimensional observational data to facilitate predictive understanding (Liu et al. 2025; Duan et al. 2022) poses a novel and interdisciplinary challenge.
@@
From Kepler’s discovery of planetary motion laws to the most advanced symbolic regression techniques, the discovery of equations has historically relied on scientists manually processing observational data to obtain compact physical quantities in suitable coordinate systems (Chen et al. 2022; Luan, Liu, and Sun 2022). However, automatically discovering both physical coordinates and the governing equations on them from high-dimensional data is a challenging problem. With advancements in deep learning, especially visual models (Krizhevsky, Sutskever, and Hinton 2012; He et al. 2016; Dosovitskiy 2020), discovering physical laws from high-dimensional data, like videos, has become more feasible. Unsupervised motion video prediction with neural networks integrated with parameterized physical modules (e.g., neural ODE (Chen et al. 2018)) allows for implicit learning of dynamics from high-dimensional observations (Guen and Thome 2020; Wu et al. 2023; Higgins et al. 2021; Hofherr et al. 2023; Li et al. 2025a; Bounou, Ponce, and Carpentier 2021). Another approach is to embed physical engines as inductive biases into the model pipeline, compelling the neural network to infer physical quantities that satisfy the governing physical laws from video streams (Kadambi et al. 2023; Yang et al. 2022; Fotiadis et al. 2023; Kandukuri et al. 2022; Jaques, Burke, and Hospedales 2019). However, these methods separate the discovery of coordinates and equations, presuming known dynamics equations or physics engines before identifying physical quantities. Recent efforts have succeeded in jointly discovering physical coordinates and equations (Champion et al. 2019; Luan, Liu, and Sun 2022; Zhang, Liu, and Sun 2024), yet they still rely on prior knowledge of equation forms and state dimensions, limiting their generalization to unfamiliar systems.
@@
High-dimensional data challenges traditional optimization methods like Bayesian optimization due to the lack of well-defined quantitative evaluations for intrinsic coordinates and governing equations. Large language models (LLMs) excel in problem-solving by leveraging pretrained knowledge and enhancing prompts for scientific discovery across various domains. (Boiko et al. 2023; M. Bran et al. 2024; Kang and Kim 2024). Multimodal large language models (MLLMs) with visual encoders achieve spatial understanding akin to human visual intelligence. (Wu et al. 2025). Recent studies (Liu et al. 2025) show MLLMs can deduce physical properties and motion patterns from images without prior knowledge, revealing their potential in high-dimensional data. Moreover, most symbolic regression methods, based on genetic programming, search equation spaces through mutation and recombination. By leveraging the pretraining knowledge of LLM to guide this search, equation discovery can be more efficient and avoid suboptimal solutions (Shojaee et al. 2024; Ma et al. 2024). Thus, we can anticipate that a well-designed reasoning framework and visual prompts could harness MLLMs for visual perception and scientific discovery, enabling automatic identification of joint physical coordinates and governing equations.
@@
Problem Fomulation
In the task of inferring governing equations for high-dimensional data, the goal is to find a compact and accurate symbolic expression. Here, we consider a class of video-like systems governed by low-dimensional coordinates, that is, systems whose dynamics evolve on a low-dimensional manifold (Champion et al. 2019). Given the dataset $\mathcal{D}=\{x_{i}\}^{n}_{i=1}$, where $x_{i}\in\mathbb{R}^{D}$ is the $i$-th observed state (i.e., a frame of video), we aim to uncover the underlying mathematical relationships such that $\frac{dx}{dt}=f(x)$. This fundamentally relies on an effective coordinate system $\mathcal{Z}$, in which the dynamics have a simple low-dimensional representation $\frac{dz}{dt}=\hat{f}(z)$. Therefore, this type of problem consists of two coupled subproblems:
@@
Discovering physical coordinates
We consider two types of high-dimensional systems. The first type is the pixel coordinate system, where physical laws are presented in the form of Cartesian coordinates of rigidly moving objects (Luan, Liu, and Sun 2022; Zhang, Liu, and Sun 2024; Liu et al. 2025). MLLM needs to identify the intrinsic physical coordinate system by locating the trajectory of the target object’s motion through a sequence of coordinates. The second type is the more complex latent coordinate system, whose spatiotemporal evolution is governed by latent low-dimensional variables, presenting periodic or quasiperiodic oscillatory patterns (Champion et al. 2019; Rudy et al. 2017). We sample a small number of frames from the data and ask MLLM to determine their type in order to initiate different locating tools.
Detection for Pixel Coordinate
@@
Plot Replayer displays MLLM’s initial locating results as markers (e.g., red dots) on the original frame and calls MLLM to compare the visual discrepancy between the target object and the detection location for error correction. As a feedback iteration step in the reasoning chain, the replayer is used after the initial locating of each frame. However, for the sake of simplicity in the formula, we have incorporated it into the $\mathrm{LocateTool}$ functionality in Algorithm 1 in appendix.
After serially detecting the state of each frame, we introduce a filter tool to smooth the physical coordinate trajectory in order to enhance the continuity of motion between frames. Specifically, we designed a feedback correction module, $\mathrm{Filter}$, based on the Savitzky-Golay filter (Savitzky and Golay 1964). We plot the recognized trajectory as a visual prompt and ask MLLM to determine the window length $h$ and polynomial order $p$ based on the evolution trend and periodicity of the trajectory. Each filtering result is visualized alongside the originally recognized trajectory as the visual prompt for the next iteration, until MLLM determines that the smoothing has preserved the evolution trend while maintaining inter-frame coordinate continuity. The pseudocode for serially extracting the physical coordinate sequence from the high-dimensional motion video is shown in Algorithm 1 in appendix.
Detection for Latent Coordinate
The task of discovering physical coordinates in the latent coordinate system is no longer visual target localization, but rather the encoding of intrinsic variables that capture complex spatiotemporal evolution patterns (Chen et al. 2022). Here, intrinsic variables refer to the minimal set of variables that describe all the information of a dynamic system (Li, Wang, and Li 2023). Although prior work has developed manifold learning methods such as Koopman operator theory and delayed embedding (Lusch, Kutz, and Brunton 2018; Wu et al. 2024), these models heavily rely on prior assumptions about the system’s intrinsic dimensions (i.e., the number of intrinsic variables).
@@
Reasoning governing equations
We adopt the sparse identification method for nonlinear dynamics (SINDy) (Brunton, Proctor, and Kutz 2016) as the starting point for our approach. Inspired by the fact that latent governing equations typically contain only a finite number of terms, SINDy identifies the equations by optimizing the coefficients of predefined symbolic terms (e.g. $sin$, $exp$, and $x^{2}$). Specifically, SINDy defines a symbolic library $\Theta(z)=[\theta_{1}(z),\theta_{2}(z),...,\theta_{k}(z)]\in\mathbb{R}^{d\times k}$, where $\theta_{i}$ represents a candidate symbolic term. The symbolic library is linearly combined to form a deterministic system $\frac{dz}{dt}=\Theta(z)\Xi$, where $\Xi=[\xi_{1},\xi_{2},...,\xi_{k}]\in\mathbb{R}^{k\times 1}$ is the coefficient to be optimized for each symbolic term. Given the coordinate sequence $Z=\{z_{i}\}^{n}_{i=1}$ and its derivative sequence $\dot{Z}=\{\dot{z_{i}}\}^{n}_{i=1}$, SINDy minimizes the loss function
@@
Hypothesis generation
Previous SINDy-based works have succeeded (Chen, Liu, and Sun 2021; Gao and Yan 2022), but rely on a high-quality predefined symbolic library, often needing prior system knowledge to manually filter out unnecessary terms. To address systems without prior knowledge, we design a hypothesis generation module to mine the physical knowledge pre-learned by MLLM and automatically select potential symbolic terms. The hypothesis generation step utilizes the frozen MLLM to propose diverse and promising candidate symbolic terms. At each iteration $t$, MLLM uses replay samples from the experience pool (Sec. Experience management) as text and visual prompts, suggesting the candidate symbolic library $\Theta_{t}$.
Hypothesis Assessment
@@
$\begin{split}L_{AE-SINDy}={}&\frac{1}{n}\sum^{n}_{i=1}||x_{i}-\psi_{d}(x_{i})||^{2}_{2}\\ &+||\nabla_{x_{i}}(z_{i})\dot{x}_{i}-\Theta_{t}(z_{i})\Xi_{t}||^{2}_{2}\\ &+\eta|\Xi|\end{split}$ (2)
@@
We employ the $R^{2}$ metric as a quantitative measure of how well the equation fits the data. Additionally, we evaluate the number $length_{t}$ of symbolic terms in the equation to quantify its simplicity. In symbolic regression tasks, the goal of maintaining equation simplicity is to avoid overfitting to the limited observational data. At this stage, the candidate symbolic library hypothesis $\Theta_{t}$ is assessed based on both the $R^{2}$ metric and the equation length to determine its $fitness_{t}$. Thus, the equation discovery process is divided into two steps: (i) constructing the symbolic library, and (ii) optimizing based on stochastic gradient using PyTorch (Kingma 2014). This process aligns with the philosophy of human scientific discovery (Ma et al. 2024; Guo et al. 2024; Huang et al. 2024), where the MLLM makes decisions on the discrete components of the SINDy symbolic library based on pretrained knowledge and prior experiences in prompts, akin to the empirical hypotheses made by human scientists during scientific discovery.
Experience management
@@
Pixel coordinate system. We validate the effectiveness of the proposed method on the pixel coordinate system using several dynamical systems shown in Figure 2, including: Linear, Cubic, Circular, Van Der Pol (VDP), Glider, and Exp equations. These equations include various characteristics such as linear terms, nonlinear terms, and significant differences in time scales (Guckenheimer 1980), which together form an evaluation benchmark across different levels of difficulty. The video data is generated by simulating physical trajectories, with details provided in the Appendix A.
@@
  • Lambda–Omega (LO) equation (Champion et al. 2019) generates spiral wave structures featuring two oscillating spatial modes.
  • Brusselator (Bruss) equation (Lopez and Atzberger 2022) models the concentration dynamics in chemical reactions, where the system’s trajectory eventually reaches a limit cycle attractor.
  • Shallow-Water (Water) equation (Takamoto et al. 2022), derived from the compressible Navier-Stokes equations,describes fluid dynamics in shallow water scenarios, commonly applied in large-scale geophysical flows and tsunami simulations.
  • Real video (Schmitt et al. 2024) visualizes the formation of Kármán vortex streets as water flows past a cylindrical obstacle with a Reynolds number of 171, using dye to highlight the vortex pattern.
@@
Experimental Setup
Models. We utilize the GPT-4o (Islam and Moushi 2024) in our experiments to explore the performance of our proposed VER. We validate the improvements of VER over the basic GPT-4o in the ablation study. The penalty coefficient $\eta$ in Equation 1 is set to 0.010.010.01 and we explore the performance of hyperparameter search in the appendix B.
Baselines. Related works (Luan, Liu, and Sun 2022; Udrescu and Tegmark 2021) on discovering coordinates from pixel coordinate videos typically employ a combination of visual localization models and symbolic regression models. We adopt a recent work (Luan, Liu, and Sun 2022), Video-SINDy, as the baseline method for the pixel coordinate system. In addition, we also replace the coordinate encoder in Video-SINDy with SAM (Kirillov et al. 2023) as a zero-shot baseline, SAM-SINDy. For the latent coordinate system, the most relevant work (Champion et al. 2019) constructs an end-to-end model combining an autoencoder and SINDy (AE-SINDy) to interpret low-dimensional patterns in high-dimensional data. Since it requires prior knowledge of the low-dimensional space dimensions, we ensure that its dimensions are consistent with those discovered by VER.
@@
System with Pixel Coordinate
@@
Figure 2 shows VER’s reasoning results in the pixel coordinate system, including governing equations and inferred trajectories compared to the optimal baseline. VER uses only raw video input, without needing motion type or equation terms. VER’s trajectories closely align with ground truth over time, demonstrating MLLM’s spatial understanding through locating prompts, as supported by (Wu et al. 2025). The inferred ODE terms perfectly match the ground truth, indicating VER’s ability to uncover intrinsic dynamics in pixel space.
@@
We assess if VER and baseline methods’ discovered equations include all correct terms, count incorrect terms, and evaluate fitting accuracy ($R^{2}$). We numerically integrate discovered equations to predict trajectories, comparing them to real ones in Table 1. VER consistently outperforms baselines, especially in long-term predictions, where inaccuracies are magnified. This highlights VER’s superior efficiency and accuracy over traditional visual symbolic regression methods. Additionally, we collect physical motion videos from public sources to confirm VER’s reliability on real systems (Appendix B).
Method Equation Prediction
Terms
Found
False
Positives
$R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear
Video-SINDy Yes 0.91±0.340.91$\pm$ 0.340.91 ± 0.34 0.95±0.020.95$\pm$ 0.020.95 ± 0.02 0.90±0.030.90$\pm$ 0.030.90 ± 0.03 0.87±0.060.87$\pm$ 0.060.87 ± 0.06
SAM-SINDy Yes 0.45±0.130.45$\pm$ 0.130.45 ± 0.13 0.90±0.030.90$\pm$ 0.030.90 ± 0.03 0.87±0.030.87$\pm$ 0.030.87 ± 0.03 0.35±0.130.35$\pm$ 0.130.35 ± 0.13
VER Yes $\mathbf{0.38\pm 0.22}$ $\mathbf{0.96\pm 0.03}$ $\mathbf{0.97\pm 0.02}$ $\mathbf{0.96\pm 0.02}$
Cubic
Video-SINDy No 3.52±1.123.52$\pm$ 1.123.52 ± 1.12 0.82±0.070.82$\pm$ 0.070.82 ± 0.07 0.68±0.180.68$\pm$ 0.180.68 ± 0.18 0.26±0.190.26$\pm$ 0.190.26 ± 0.19
SAM-SINDy No 2.17±1.202.17$\pm$ 1.202.17 ± 1.20 0.76±0.090.76$\pm$ 0.090.76 ± 0.09 0.70±0.120.70$\pm$ 0.120.70 ± 0.12 0.22±0.180.22$\pm$ 0.180.22 ± 0.18
VER Yes $\mathbf{1.41\pm 0.73}$ $\mathbf{0.95\pm 0.03}$ $\mathbf{0.80\pm 0.16}$ $\mathbf{0.62\pm 0.28}$
Circular
Video-SINDy Yes 0.31±0.100.31$\pm$ 0.100.31 ± 0.10 0.98±0.010.98$\pm$ 0.010.98 ± 0.01 0.98±0.030.98$\pm$ 0.030.98 ± 0.03 0.90±0.100.90$\pm$ 0.100.90 ± 0.10
SAM-SINDy Yes 0.12±0.030.12$\pm$ 0.030.12 ± 0.03 0.92±0.010.92$\pm$ 0.010.92 ± 0.01 0.99±0.010.99$\pm$ 0.010.99 ± 0.01 0.93±0.040.93$\pm$ 0.040.93 ± 0.04
VER Yes 0 $\mathbf{0.99\pm 0.01}$ $\mathbf{1.00\pm 0.00}$ $\mathbf{0.97\pm 0.05}$
VDP
Video-SINDy Yes 2.31±0.652.31$\pm$ 0.652.31 ± 0.65 0.89±0.080.89$\pm$ 0.080.89 ± 0.08 0.82±0.060.82$\pm$ 0.060.82 ± 0.06 0.49±0.130.49$\pm$ 0.130.49 ± 0.13
SAM-SINDy Yes 2.45±0.592.45$\pm$ 0.592.45 ± 0.59 0.88±0.030.88$\pm$ 0.030.88 ± 0.03 0.92±0.040.92$\pm$ 0.040.92 ± 0.04 0.67±0.080.67$\pm$ 0.080.67 ± 0.08
VER Yes $\mathbf{0.80\pm 0.40}$ $\mathbf{0.97\pm 0.01}$ $\mathbf{0.94\pm 0.02}$ $\mathbf{0.73\pm 0.07}$
EXP
Video-SINDy Yes 2.76±0.882.76$\pm$ 0.882.76 ± 0.88 0.93±0.080.93$\pm$ 0.080.93 ± 0.08 0.93±0.020.93$\pm$ 0.020.93 ± 0.02 0.90±0.030.90$\pm$ 0.030.90 ± 0.03
SAM-SINDy Yes 3.21±1.173.21$\pm$ 1.173.21 ± 1.17 0.95±0.020.95$\pm$ 0.020.95 ± 0.02 0.96±0.010.96$\pm$ 0.010.96 ± 0.01 0.95±0.020.95$\pm$ 0.020.95 ± 0.02
VER Yes $\mathbf{1.34\pm 0.29}$ $\mathbf{0.99\pm 0.01}$ $\mathbf{0.99\pm 0.01}$ $\mathbf{0.98\pm 0.02}$
Glider
Video-SINDy No 2.18±0.492.18$\pm$ 0.492.18 ± 0.49 0.94±0.020.94$\pm$ 0.020.94 ± 0.02 0.98±0.010.98$\pm$ 0.010.98 ± 0.01 0.91±0.030.91$\pm$ 0.030.91 ± 0.03
SAM-SINDy No 2.23±0.562.23$\pm$ 0.562.23 ± 0.56 0.85±0.040.85$\pm$ 0.040.85 ± 0.04 0.99±0.040.99$\pm$ 0.040.99 ± 0.04 0.32±0.270.32$\pm$ 0.270.32 ± 0.27
VER Yes $\mathbf{0.84\pm 0.17}$ $\mathbf{0.98\pm 0.01}$ $\mathbf{0.99\pm 0.01}$ $\mathbf{0.92\pm 0.02}$
@@
Table 1: Average performance of pixel coordinate systems over 10 runs with varying seeds. $R^{2}@n$ indicates $R^{2}$ for $n$ steps. Best results are in bold.